### v1

# A General Bayesian Framework for Informative Input Design in System Identification

Abstract
We tackle the problem of informative input design for system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters. Our algorithm consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and propose an approximate iterative method to optimize the model’s parameters. Based on this Bayesian formulation, we are able to define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. Our method outperforms model-free baselines with various linear and nonlinear dynamics.
I Introduction
Controlling uncertainty is a central problem for various engineering fields. It abstracts to the selection of inputs to a system in order to reduce an uncertainty measure for a variable of interest. In robotics, the problem usually coincides with adaptive informative input design, where trajectories are planned to maximize the information gathered from the unknown environment [1]. Similarly, robots may choose actions to reduce the uncertainty in their state [2]. In systems engineering, we may want to choose inputs to a system to improve the estimate of system dynamics [3].
In this work, we are interested in the problem of informative input design for the modeling of a system. Our formulation tackles the general problem of dataset collection for model-fitting purposes. Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [4]. Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging. Informative input design can also be regarded as the first step in a more complicated pipeline that involves the control of the underlying system, e.g., with model predictive control [5]. The problem of selecting input sequences that lead to highly informative measurements under constraints is still an open problem [4].
There are four basic elements in the problem of informative input design. We need to select a parametric family of models for the underlying system. Our parameter estimate is improved once more data becomes available and we must choose the optimization approach to update the model parameters. The model family must be expressive enough to capture the true system, but also simple enough to allow the optimization algorithm to converge in a meaningful manner. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. The last minimization problem must also be accompanied by an optimization approach.
These specifications make designing methods for informative input design challenging and lead to algorithms that are usually limited to a specific set of parametric models. Our approach allows for the selection of any parametric family of models for the true system, although we do not focus on the suitability of the chosen model. It leverages first-order information for the model and successive convex problems based on the first-order Taylor expansion to update the model parameters. This simplifies the model update step discussed in the previous paragraph. Because the first-order Taylor expansion gives rise to a linear Gaussian model, we use an uncertainty measure that is accurate in the linear Gaussian case for the input design.
The rest of the paper is organized as follows. In section II we discuss prior work. In section III, we formulate the problem of informative input design, while in section IV we describe our proposed approach. We test our algorithm in various scenarios in section V and the paper concludes with section VI.
II Prior Work
The problem of informative input design has been well studied. Early research on the optimal selection of inputs dates back to the 1960s [6, 7] and highlights the importance of the problem for the identification of systems. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [8, 9]. This section summarizes three categories of informative input design: model-free methods, methods for linear systems, and methods for nonlinear systems.
Informative inputs can be designed by ignoring the current estimate of the system dynamics. These methods are called model-free because they are not based on a model of the system. One such method is the orthogonal optimized multisine input design [10], which has found applications in the study of aircraft stability and control. This technique produces inputs that are mutually orthogonal in both the time and frequency domains, in an attempt to perturb the system in orthogonal directions. Similarly, pseudorandom binary sequences can be used to excite the system [11]. The idea is to leverage the sequence’s randomness in order to excite the system in interesting ways. Our method differs from model-free approaches because it uses the currently estimated model to inform the next input design.
Although they are simple, linear models are a powerful representation tool. They usually allow for closed-form analysis and are easily interpretable from a statistical standpoint. Informative input design using such models has a vast literature. A recent method that aims to estimate linear dynamics suggests selecting the next input by minimizing the trace of the covariance of the model parameters with linear dynamics constraints [4]. By linearizing the objective, the authors obtain a convex problem that is easily solvable for the next input. This method is compatible with the dynamic mode decomposition with control (DMDc) framework that produces reduced order models for the dynamics [12]. Our method can provide an approximate solution to any parametric model family and is exact in the case of linear models.
In various applications, the dynamics contain modes that are not properly described by a linear model. In these cases, we must design inputs for nonlinear system identification. Early results focused on specific dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification. In this case, convex-based methods exist for informative input design [15]. In contrast with these methods, our proposed algorithm is applicable in the case of general nonlinear models.
Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16]. Our work adds to this work in an important way: we present a method to update the estimate of the model parameters that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [17], who further propose an extended Kalman filter for the parameter estimation. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input. Our objective depends on a parameter that directly encourages the satisfaction of any constraints, i.e., dynamics constraints. In contrast with both works [16, 17], we present a fine-tuning method for the assumed noise covariances that can improve convergence.
III Problem Formulation
We assume that a system $g$ is characterized by the relation
$y=g(x),$ (1)
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We refer to $x$ as the input and $y$ as the output.
We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{q}$. At every iteration, our goal is to compute:
  1. 1.
    $\hat{\theta}$ such that $g(\cdot)\approx f(\cdot;\hat{\theta})$ using the available data,
  2. 2.
    the optimal next input to the system that decreases our uncertainty over the unknown parameter $\theta$.
We opt for a Bayesian treatment of the problem. We suppose that the inputs and outputs are exactly observed, because our main goal is efficient dataset acquisition. Our method easily extends to the case of noisy observation of the outputs.
IV Proposed Method
We present our methodology for the design of an informative input in system identification. We assume that we have already observed input-output data points $(x_{i},y_{i})_{i=1}^{n}$ from the dynamics:
$y_{i}=g(x_{i}).$ (2)
We first outline our Bayesian approach for improving our current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. Finally, we discuss some fine-tuning that can improve performance before bringing the concepts together in our proposed algorithm.
IV-A Parameter Estimation
The collected data satisfy eq. 2. Because the current model $f(\cdot;\theta)$ might not match the true system $g(\cdot)$ exactly, it holds that
$y_{i}=f(x_{i};\theta)+\epsilon_{i},$ (3)
for some noise realization $\epsilon_{i}$. We choose the underlying noise model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma_{i})$. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
$p(\theta\mid y_{1},\dots,y_{n})\propto p(\theta)\prod_{i=1}^{n}p(y_{i}\mid\theta),$ (4)
where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$. The noise and prior are assumed Gaussian because they lead to a convex maximum likelihood problem [18].
The maximum a posteriori estimate for the parameter under these assumptions is given by
$\hat{\theta}_{+}$ $=\operatorname*{arg\;max}_{\theta}\log p(\theta\mid y_{1},\dots,y_{n})$ (5)
$=\operatorname*{arg\;min}_{\theta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}+$
$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma_{i}^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver.
Suppose that the current parameter estimate is $\hat{\theta}$. We do not allow our next estimate to differ much relative to the current one. This is justified because the next input-output pair should not change our estimate drastically. For model parameters close to $\hat{\theta}$, the parametric family can be accurately approximated using a first order Taylor series. We write
$f(x;\theta)$ $\approx f(x;\hat{\theta})+\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\left(\theta-\hat{\theta}\right)$ (6)
$=b(x;\hat{\theta})+C(x;\hat{\theta})\theta$
for $\theta$ close to $\hat{\theta}$, where
$\left[\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\right]_{i,j}$ $=\dfrac{\partial f_{i}}{\partial\theta_{j}}|_{(x;\hat{\theta})},$ (7)
$b(x;\hat{\theta})$ $=f(x;\hat{\theta})-\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\hat{\theta},$
$C(x;\hat{\theta})$ $=\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}.$
Plugging (6) into (5) and imposing the distance constraint, we get that
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta\lVert\hat{\theta}\rVert}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x;\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma_{i}^{-1}}^{2}.$
This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19]. The constraint guarantees that we search for the next model parameter in the region where the current Taylor series is accurate. It corresponds to a trust region.
IV-B Informative Input Design
After obtaining $\hat{\theta}_{+}$, our goal is to select $x_{n+1}$. The goal of selecting $x_{n+1}$ is to reduce our uncertainty in the posterior estimate of $\theta$. Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under completely linear Gaussian assumptions, the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
It has been shown that such an expression can be accurate in the nonlinear case, although it is exact only in the linear regime [20].
To minimize the posterior covariance, we maximize some notion of magnitude,
$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
of the posterior inverse covariance over $x_{n+1}$. We can do so because the uncertainty in the parameter estimate is proportional to the magnitude of the posterior covariance, which is inversely proportional to the magnitude of the inverse posterior covariance. For example, we can use $\log\det$ as $\mathcal{M}$. The determinant of the covariance equals the product of its eigenvalues, which are proportionally related to the semi-axes of the probability contours for a Gaussian distribution.
In the process of choosing $x_{n+1}$, there usually exist constraints. For example, the input might need to belong to a specific set, $x_{n+1}\in\mathcal{X}$. The same might hold for the corresponding output, $g(x_{n+1})\in\mathcal{Y}$. We can take such constraints into account by solving the following optimization problem:
$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
$+\lambda\left(\mathbbm{1}(x_{n+1}\in\mathcal{X})+\mathbbm{1}(f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y})\right).$
In order to solve eq. 11 with gradient descent methods, we can replace the indicators with smooth functions. By increasing parameter $\lambda\geq 0$, we can move the search towards feasible inputs $x_{n+1}$. This is an example of a penalty method [21]. As a receding horizon extension, we can easily extend eq. 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$.
Equation 9 defines the posterior covariance for the parameter $\theta$. At the next iteration , we can replace the constraint of eq. 8 with
$\lVert\theta-\hat{\theta}\rVert^{2}_{\Sigma_{\mathrm{posterior}}^{-1}}\leq\chi_{d,p}^{2},$
where $\chi_{d,p}^{2}$ is the value of the chi-squared distribution of $d$ degrees of freedom at the $p$-percentile. This would allow $\theta$ to vary within the contours that correspond up to the $p$-th percentile of the posterior Gaussian distribution. Choosing the adequate constraint in eq. 8 involves balancing this statistical interpretation with the accuracy of the Taylor series.
IV-C Updating the Covariances $\Sigma_{i}$
At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$. After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\hat{\epsilon}_{i}$ by comparing the true output to the expected output, i.e., for $i=1,\dots,n$
$\hat{\epsilon}_{i}=y_{i}-f(x_{i};\hat{\theta}_{+}).$ (12)
Using these samples, we can compute a sample covariance for the overall error in our model
$\hat{\Sigma}_{+}=w\Sigma_{\mathrm{meas,prior}}+\dfrac{1-w}{n}\sum_{t=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top},\ w\in\left[0,1\right],$ (13)
i.e., a convex combination of the empirical covariance and a prior covariance. This can be utilized in later iterations.
IV-D Algorithm
We combine the modules presented above in algorithm 1. We assume we have already acquired dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ and show the proposed procedure for augmenting the dataset by one point.
Notice that when computing $x_{n+1}$ in eq. 11 we choose $\Sigma_{i}\leftarrow\hat{\Sigma}_{+}$ from (13) for $i=1,\dots,n+1$. When performing the maximum a posteriori estimation in eq. 8 , however, we use $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. This selection is made because our next input should excite the system in a way that is informative and decreases the uncertainty present in our current system estimate. However, when improving the estimate of the model parameters we should use less informative covariances to robustify the estimation.
Algorithm 1 Proposed Algorithm for Parameter Estimation and Input Design
1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma_{\mathrm{meas,prior}}$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta,\lambda$ and $w$
2:$\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}},\ \mathrm{for\ all}\ i=1,\dots,n$
3:Obtain updated estimate $\hat{\theta}_{+}$ by eq. 8
4:$\Sigma_{i}\leftarrow\hat{\Sigma}_{+},\ \mathrm{for\ all}\ i=1,\dots,n+1$
5:Obtain next input $x_{n+1}$ by solving eq. 11
6:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
7:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$
V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design We further perform an ablation study to show the importance of finetuning the covariances $\Sigma_{i}$ in our algorithm.
Baselines
Our emphasis is on the informative input design. We compare against three different model-free baselines for the selection of inputs. We maintain the same parameter estimation module from section IV in all methods. The first baseline is random. At every iteration, we run algorithm 1, with a change in line 4: we suppose $x_{n+1}\sim\mathcal{N}(0,I)$, i.e., $x_{n+1}$ is randomly selected. The second baseline is based on pseudorandom binary sequences and is called PRBS. At every iteration, we run algorithm 1, with a change in line 4: we suppose that each coordinate of $x_{n+1}$ originates from a pseudorandom binary sequence. Finally, the last baseline is called mutlisine and is based on orthogonal multisine inputs [10]. In line 4 of algorithm 1, we set the coordinate $i$ of $x_{n+1}$ as follows
$x_{n+1,i}=\sum_{j=0}^{M-1}A_{i,j}\sin\left(\dfrac{2\pi(d_{x}j+i)(n+1)}{T}+\phi_{i,j}\right),$
where $M$ is the total number of sinusoidal components and $T$ is the total number of inputs to be designed. We sample each phase $\phi_{i,j}$ uniformly in $\left[0,2\pi\right]$ and each amplitude $A_{i,j}$ uniformly in $\left[-1,1\right]$. We select $M=30$ in all experiments below.
Algorithm Parameters
We set $w=0.2$, $\delta=0.3$, and $M=30$. We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition. We do set an upper bound on the allowed input magnitude. If the input provided by any method has a larger magnitude, it is scaled to have the largest allowed magnitude. The first datapoint is assumed fixed for all methods.
Metrics
In all cases below , the true system belongs to the parametric model family. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq. 9. We assume $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs. The results are gathered in table I for all experiments.
V-A Linear System
We first consider a linear system
$g(x_{t})=\begin{bmatrix}1&2\\ 3&4\end{bmatrix}x_{t}$ (14)
and the parametric model
$f(x_{t};\theta)=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}x_{t}.$ (15)
Our results are shown in Figure 2. In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the various methods. Our algorithm exhibits a faster convergence to the true model parameters compared to the baselines. This is verified by the lower value of $\log\det$ for the final error covariance.
V-B Hénon map dynamics
We consider the Hénon map dynamics
$g(x_{t})=\begin{bmatrix}1-\alpha x_{t,1}^{2}+x_{t,2}\\ \beta x_{t,1}\end{bmatrix},$ (16)
where $\alpha=1.4$ and $\beta=0.3$. It is one of the most studied systems with chaotic behavior [22]. The parametric model is
$f(x_{t};\theta)=\begin{bmatrix}1-\theta_{1}x_{t,1}^{2}+x_{t,2}\\ \theta_{2}x_{t,1}\end{bmatrix}.$ (17)
Our results are shown in Figure 3. Our proposed algorithm is able to converge to the true model parameters and outperforms the baselines with respect to both metrics considered. From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter the estimate has a higher uncertainty in the second coordinate $\theta_{2}$, which is captured by the elongated ellipse axes along that direction.
V-C Unicycle Model
We also evaluate our algorithm on the unicycle dynamics model, which is nonlinear and given by the equation
$x_{t+1}=\begin{bmatrix}x_{t,1}+u_{1,t}\Delta t\cos(x_{t,3})\\ x_{t,2}+u_{1,t}\Delta t\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\Delta t\end{bmatrix},$ (18)
where $u_{1,t}$ is the linear velocity input and $u_{2,t}$ is the angular velocity input. We use $\Delta t=0.1$. In this example, we will only be able to choose the control input $u_{t}$ at every timestep. In other words, the informative input design must be done in sequence, with the next state depending on the previous state and input. The parametric model is assumed to be
$f(x_{t},u_{t};\theta)=\begin{bmatrix}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}.$ (19)
The time discretization is thus one of the unknown parameters. We start the model from the origin with zero being the first control for all methods. Figure 4 shows the performance of our algorithm. It outperforms the considered baselines. We further plot $(x_{t,1},x_{t,2})$ for the various methods.
V-D Ablation
In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with a variant that uses $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$ throughout. The new variant is termed proposed (no finetuning). In this simulation , we remove the amplitude constraint. It is evident in Figure 1 that the variant does not converge properly. This highlights the importance of finetuning the covariances $\Sigma_{i}$.
Figure 1: Ablation study in the case of Hénon dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
Figure 2: Results for the linear dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
Figure 3: Results for the Hénon dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) State of dynamics as a function of timestep for the various methods.
Figure 4: Results for the unicycle dynamics.
Method/Experiment Linear Dynamics Hénon Dynamics Unicycle Dynamics
proposed $\mathbf{-18.5}$ $\mathbf{-14.9}$ $\mathbf{-9.8}$
random $-10.7$ $-5.9$ $-7.0$
PRBS $-10.2$ $-5.1$ $-9.2$
multisine $-13.7$ $-9.7$ $-8.9$
TABLE I: $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep for the various dynamics studied.
VI Conclusion
We present a Bayesian methodology for informative input design that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariances that greatly impacts performance. Our algorithm also has limitations. We must select the trust region shape and volume when performing the parameter estimation , and we do not address the problem of selecting the parametric model family. Future work will propose a method for judging the suitability of a given parametric model family.
Acknowledgements
Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. The NASA University Leadership Initiative (grant ##\##80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity. For the first author, this work was also partially funded through the Alexander S. Onassis Foundation Scholarship program. We would like to thank Joshua Ott and Arec Jamgochian for their feedback.